### v1

@@ 1 Introduction
Using data from the US Energy Information Administration on oil production and price and data available through FRED on global real economic activity, we estimate the structural vector autoregressive model of the global crude-oil market proposed by Kilian (2009a). The updated results provide further evidence that Kilian’s (2009a) conclusions hold. In addition, a procedure that accounts for conditional heteroskedasticity of unknown form in the VAR model, indicates that Kilian’s conclusions are even stronger than initially stated (Brüggemann, Jentsch, and Trenkler, 2016).
@@
The structural VAR model disentangles supply shocks from demand shocks in the global crude-oil market. In the model, supply comprises global oil production and demand comprises two components—a component that reflects the global business cycle and a component that reflects oil-specific demand (buyers concerned about availability in the future will buy out of precaution, for example). This allows Kilian (2009a) to show how each shock separately determines fluctuations in the real price of crude oil. Moreover, how real Gross Domestic Product responds to an unanticipated increase in the price of oil depends on the nature of the shock. The same conclusion applies to the Consumer Price Index. Kilian’s results explain the macroeconomic implications of oil-price fluctuations.
@@
In a different direction, we use the updated series of structural shocks to learn about another important policy question: how does a local labor market respond to the identified structural shocks? The question is of interest to policymakers in California. California has adopted the goal of carbon neutrality by 2045 or earlier (CARB, 2022). In Kern County, which accounts for 70 percent of California’s oil production, there is concern about potential job loss and lost tax revenue as the economy transitions away from fossil fuels. Yet, how much Kern’s economy depends on oil is hard to assess. We document that the unemployment rate in Kern responds meaningfully to identified shocks in the global market for oil and that the response depends on the nature of the shock. This is suggestive evidence that policymakers may need to consider expanding options in Kern through, for example, place-based policy. [^1] What will happen in California as 2045 approaches is impossible to know with certainty. In particular, a backward-looking structural VAR model will not be a good predictor of abrupt change. Instead, we view Kern’s responses to identified structural innovations as suggestive evidence that the Kern economy depends on oil. The exercise shows how Kilian’s (2009a) pioneering methods and conclusions apply more generally.
@@
Questions about how oil affects the US economy are not fully answered. Further research will benefit from other contributions of our replication. We port Kilian’s original MATLAB code to R. Some researchers may prefer to work in R as the R ecosystem makes tidying data easier. We use the targets package to ensure replicability. The targets package provides Make-like tools that automatically make any necessary changes to dependencies when any change is registered. In addition, we retrieve data from the US Energy Information Administration and FRED using code, which makes estimating the most recent structural shocks as straightforward as rerunning the code. These resources will make addressing questions easier in the ongoing debate about the effects oil-price shocks have on the US economy. [^2] An introduction to Make is provided by Jesús Fernández-Villaverde in lecture notes titled “Computational Methods for Economists.”
@@
We present a replication of Kilian’s (2009a) results in section 2. We conclude that Kilian (2009a) should be praised for writing and sharing code that makes reproducing results straightforward (Kilian, 2009b). The replication comprises a description of the data (section 2.1), a specification of the structural VAR model (section 2.2), a presentation of the estimated impulse response functions (section 2.3), a decomposition of the real price of oil into identified components (section 2.4), and a second-stage analysis of how US real GDP and the Consumer Price Index respond to the identified structural shocks. In section 3 we strengthen Kilian’s (2009a) conclusions by estimating the structural VAR model using data from February 1974 to March 2024 and constructing confidence intervals for impulse responses that are robust to arbitrary forms of conditional heteroskedasticity. To demonstrate the wider applicability of the results, in section 4 we analyze how the global oil market affects outcomes in a local labor market.
@@ 2 Replication
2.1 Data
Separating supply shocks from demand shocks requires data on oil production and the price of oil. Data on production are available from the US Energy Information Administration (EIA). [^3] Using the link in the text, the data can be found under “Petroleum and Other Liquids” and “Crude oil including lease condensate.” Data on price are also available from the EIA under the series U.S. Crude Oil Composite Acquisition Cost by Refiners. The EIA website includes links for retrieving the data through a web application programming interface or API. Through the API, it is possible to send structured HTTP requests that return data with JSON format or JavaScript Object Notation format. The code that gathers our data uses the EIA’s API to retrieve the latest data for production and price.
The time series that measures the global business cycle is the Index of Global Real Economic Activity, which is available through FRED. [^4] Federal Reserve Bank of Dallas, Index of Global Real Economic Activity [IGREA], retrieved from FRED, Federal Reserve Bank of St. Louis; https://fred.stlouisfed.org/series/IGREA. The series “is derived from a panel of dollar-denominated global bulk dry cargo shipping rates and may be viewed as a proxy for the volume of shipping in global industrial commodity markets.” Kilian (2009a) initially proposed this measure of the global business cycle. Kilian and Zhou (2018, 55) discuss the merits of the series, concluding “that the Kilian (2009a) index and indicies based on real commodity prices have clear advantages in modeling industrial commodity markets compared with alternative indicators of global real economic activity, including proxies for global industrial production.”
Figure 1 depicts the three series used in the structural VAR model. All series are monthly. The top panel depicts the change in global crude-oil production. The solid, blue line shows the data used by Kilian (2009a), and the dashed, red line shows data retrieved using the code for this paper (see Kilian, 2009b, for the original data and code). The original series begin in February 1973 and go through December 2007. The updated series begin in February 1974 and go through March 2024. The discrepancy occurs because data on production begin in 1974. Kilian (2009a) extends the series through 1973 using a procedure described by Barsky and Kilian (2001), a procedure that is not undertaken here. The two series largely coincide. [^5] As can be confirmed in the code, the original production series was multiplied by 10 in order to produce the plot. The reason for this transformation is unclear. Scaling any series will have no meaningful effect on the structural VAR model. Nevertheless, it speaks to the aims of the Journal of Comments and Replications in Economics, which emphasize code that reproduces results from original data sources.
Figure 1: Determinants of the global market for crude oil
Note: Top, change in global crude-oil production; middle, global real economic activity measured using the index proposed by Kilian (2009a); bottom, real price of oil measured by the acquisition cost of refiners and deflated using the Consumer Price Index for All Urban Consumers. The solid, blue lines show the data originally used by Kilian (2009a). The dashed, red lines show the updated data. All data are monthly. The updated data begin in February 1974 and go through March 2024. The original data begin one year earlier.
Sources: US Energy Information Administration, Federal Reserve Bank of Dallas, and US Bureau of Labor Statistics
The middle panel in figure 1 shows the time series of global real economic activity. Again, the solid, blue line depicts the original data and the dashed, red line depicts the updated data. A few years before 2005, the two series diverge meaningfully. This discrepancy has to do with a coding error in the original analysis described by Kilian (2019).
The bottom panel in figure 1 shows the time series for the real price of crude oil. Here the difference between the original and updated series is unknown to us. Nevertheless, the correlation between the two series is 0.98.
As Kilian (2009a) demonstrates, the three time series in figure 1 are key determinants of the global market for crude oil.
@@ 2.2 The structural VAR model
To be precise about how the determinants of the global market for crude oil are related, consider a VAR model for $y_{t}=\left(\Delta\text{oil production},\text{real activity},\text{real price of oil}\right)^{\prime}$. These are the three series shown in figure 1. The structural VAR representation is
@@
where the $B_{i}$ are $3\times 3$ parameter matricies and $\varepsilon_{t}\sim\left(0,\Sigma_{\varepsilon}\right)$; that is, $\varepsilon_{t}$ denotes a vector of serially and mutually uncorrelated innovations. The covariance matrix $\Sigma_{\varepsilon}$ is diagonal. The corresponding reduced-form model is
@@
Within the system of three equations, the first equation indicates that the stochastic short-run supply curve is vertical. In other words, oil production does not respond to innovations in demand within the month. The maintained assumption is that decisions about oil supply are based on medium-run expectations and, because adjusting supply is costly, short-run demand innovations are postulated to have a negligible affect on medium-run expectations. Evidently, the model allows energy prices to respond to all past information. There are two demand shocks and each is identified by delay restrictions. The short-run demand curve is downward sloping and can be shifted by innovations to aggregate demand and innovations to oil-specific demand (such as precautionary-demand innovations). The delay restriction is that oil-specific demand shocks do not affect global real economic activity as measured by fluctuations in shipping rates relative to trend. These postulations are further discussed by Kilian (2008) and Kilian (2009a).
@@
Inference is based on the wild bootstrap discussed by Gonçalves and Kilian (2004) with 1,000 replications. Kilian (2009a) uses 2,000 replications but the results are similar. The wild bootstrap allows for conditional heteroskedasticity of unknown form. But the asymptotic validity of the wild bootstrap is for the reduced-form impulse responses and not the asymptotic distribution of the error–covariance matrix. This feature is shown by Brüggemann, Jentsch, and Trenkler (2016) and discussed by Kilian and Lütkepohl (2017). In section 3, we adopt the inference procedures proposed by Brüggemann, Jentsch, and Trenkler (2016) to investigate whether the larger confidence intervals matter in practice.
@@
The structural impulse responses define the responses of elements of $y_{t}$ to one-time impulses in $\varepsilon_{t}$. Figure 2 shows how global oil production, real economic activity, and the real price of oil respond to one-standard deviation structural innovations. The innovations are normalized so that each innovation will tend to raise the price of oil or, put differently and slightly imprecisely, figure 2 shows negative supply shocks and positive demand shocks. Figure 2 is comparable to figure 3 in Kilian (2009a).
@@
Note: Solid lines show the point estimates and the shaded regions show two-standard-error bands. Estimates are based on the structural VAR model in (1), using the original data (Kilian, 2009b) (February 1973–December 2007). Confidence intervals were constructed using a recursive-design wild bootstrap. Dotted lines show the point estimates and two-standard-error bands produced by Kilian’s (2009a) code. The striped effect indicates the two sets of point estimates agree.
Figure 2 uses data provided by Kilian (2009b). These are the data shown in figure 1 with solid, blue lines. The data start in February 1973 and end in December 2007. The goal of figure 2 is to assess whether the R code written for the replication yields similar results to the MATLAB code that produces the original graphics.
Figure 2 confirms that the two sets of code produce similar results. The solid, blue lines show the point estimates of the structural impulse responses. They match those produced by the code provided by Kilian (2009a) exactly. Kilian’s (2009a) original estimates are shown using red, dotted lines. These overlay the solid, blue lines, which produces a striped effect. The shaded areas depict the two-standard-error bands constructed by the wild bootstrap. The dotted, red lines show those produced by the original code. Again, there is overwhelming overlap, which gives confidence that the two sets of codes are working.
Figure 3 uses our code base and our data over the original sample period. These data are the dashed, red lines in figure 1. There are two differences to keep in mind when comparing figure 2 to figure 3. First, the updated data on production start in February 1974. Second, the data on global real economic activity have been corrected (Kilian, 2019). Notwithstanding these differences, the patterns are strikingly similar.
@@
An unexpected disruption to oil supply causes an immediate decline in oil production. The decline is partially reversed within the first 5 months, but “this pattern is consistent with the view that oil supply contractions in one region tend to trigger production increases elsewhere in the world” (Kilian, 2009a, 1061–1062). The fall in production causes a rise in the price of oil, but the rise is not overwhelmingly statistically significant. In figure 3, the disruption has little effect on real economic activity, which differs slightly from Kilian’s (2009a) observation that a supply disruption “causes a small temporary reduction of real economic activity that is partially statistically significant” (1,062). (The assessment is based on one-standard-error bands, which are not shown here. Figure 3, constructed with all available data, shows one-standard-error bands.)
Turning to the second row of figure 3, the effect of an unanticipated rise in aggregate demand causes oil production to rise about 6 months after the shock. The expansion in oil production lasts until 15 months after the shock. The unanticipated rise in aggregate demand causes a significant rise in global real economic activity. The effect is persistent. Expansions in real economic activity cause a significant rise in the real price of oil. The real price of oil rises on impact (although not statistically significantly) and the rise continues for a year.
The third row of figure 3 shows the effect of an unanticipated rise in oil-specific demand. An unanticipated rise causes oil production to briefly fall, but over the 15-month horizon, oil-specific-demand increases do not cause oil production to change much. Real economic activity in response to the unanticipated rise in oil-specific demand expands—albeit modestly, if at all. What is most clear, though, is that the unanticipated rise in oil-specific demand causes the real price of oil to jump upward immediately and continue to rise for a few months, which is consistent with models of precautionary demand. The rise in the real price of oil is highly statistically significant.
The results in 3, which use the replication code and updated data, are entirely consistent with the results in Kilian (2009a).
@@ 2.4 The cumulative effect of oil demand and oil supply shocks on the real price of oil
How much each structural shock explains of the historically observed price of crude oil can be assessed through a historical decomposition of the covariance-stationary VAR model. The historical decomposition decomposes $y_{t}$ into structural innovations through its moving-average representation. It attributes the cumulative effect of the structural innovations in oil production, real economic activity, and oil-specific demand to the real price of oil at each point in time.
Figure 4 shows the cumulative contribution of oil-supply, aggregate-demand, and oil-specific-demand shocks to the real price of oil in separate panels. Each panel shows 4 lines. The top panel shows the cumulative effect of oil-supply shocks. The solid, purple line is generated from the original code and original data. The dashed, beige line is generated by using the replication code together with the original data. The two series are numerically equal, which can be seen by the striped pattern. The dot–dashed, orange line is generated by using the replication code and the updated data through December 2007. The dotted, pink line is generated by using the replication code and updated data through March 2024.
@@
Note: Top, how much oil-supply shocks explain of the historically observed fluctuations in the real price of crude oil; middle, how much aggregate-demand shocks explain of the historically observed fluctuations in the real price of crude oil; bottom, how much oil-specific-demand shocks explain of the historically observed fluctuations in the real price of crude oil. Estimates are based on the structural VAR model in (1). The solid, purple line is generated by the original code and data. The dashed, beige line, which lies on top of the purple line, is generated by the replication code and the original data. The dot–dashed, orange line is generated by the replication code using data from February 1974 to December 2007. The dotted, pink line is generated by the replication code using data through March 2024.
@@
The middle panel of 4 shows that aggregate-demand shocks cause prolonged and persistent changes in the price of oil. The bottom panel of 4 shows that shocks to oil-specific demand are responsible for the fluctuations in the price of oil. As Kilian (2009a) points out, this statistical understanding of crude-oil prices is inconsistent with narrative hypotheses about supply disruptions in the Middle East. There may be supply disruptions, but the disruptions affect firms that purchase oil, fearing further supply disruptions. These precautionary-demand motives cause the price of oil to fluctuate.
Figure 4 is directly comparable to Kilian’s (2009a) figure 4. The conclusions drawn from the figure largely agree—except for the cumulative effect the aggregate-demand shock has on the real price of crude oil using the updated series through December 2007 (the middle panel). The discrepancy can be explained by the updated series for real global economic activity and the real price of oil in figure 1. The end of the updated series for real activity is characterized by a sharp rise compared to the original series. The beginning of the updated series for the price of oil is lower than the original series. Before 1990 in the updated data, a given level of real economic activity is associated with a lower real price of oil compared to the original data. After 1990 it will not be surprising then that the estimated model using updated data predicts that the real price of oil will be relatively low, producing a negative error, which shows up in the historical decomposition. Comparing the original data to the updated replication series in figure 1 through 2007, the beginning of the series for the real price of oil and the end of the series for real economic activity are challenging for the model to match.
The discrepancy, however, is entirely absent when the structural model is estimated using data through March 2024. Therefore, going forward results will be shown for estimates produced using the full sample.
2.5 The demand and supply shocks in the global oil market affect the US economy in different ways

### v2

@@ 1 Introduction
Using data on oil production and price from the US Energy Information Administration and data on global real economic activity available through FRED, we estimate the structural vector autoregressive model of the global crude-oil market proposed by Kilian (2009a). The updated results provide further evidence that Kilian’s (2009a) conclusions hold. In addition, a procedure that accounts for conditional heteroskedasticity of unknown form in the VAR model, indicates that Kilian’s conclusions are even stronger than initially stated (Brüggemann, Jentsch, and Trenkler, 2016; Kilian and Lütkepohl, 2017).
@@
The structural VAR model disentangles supply shocks from demand shocks in the global crude-oil market. In the model, supply comprises global oil production and demand comprises two components: (i) a component that reflects the global business cycle and (ii) a component that reflects oil-specific demand (buyers concerned about availability in the future will buy out of precaution, for example [Alquist and Kilian, 2010]). This allows Kilian (2009a) to show how each shock separately determines fluctuations in the real price of crude oil. Moreover, how real Gross Domestic Product responds to an unanticipated increase in the price of oil depends on the nature of the shock. The same conclusion applies to the Consumer Price Index. Kilian’s results explain the macroeconomic implications of oil-price fluctuations.
@@
In a different direction, we use the updated series of structural shocks to learn about another important policy question: how does a local labor market respond to these shocks? The question is of interest to policymakers in California. California has adopted the goal of carbon neutrality by 2045 or earlier (CARB, 2022). In Kern County, which accounts for 70 percent of California’s oil production, there is concern about potential job loss and lost tax revenue as the economy transitions away from fossil fuels. Yet, given the complex linkages between oil extraction and the rest of the economy, Kern’s dependence on oil is hard to assess (Golding and Kilian, 2022, describe one complex link). We document how Kern’s labor market responds to identified shocks in the global market for oil. And we find that the response depends on the nature of the shock. Kern’s dependence on oil suggests that economic forces will restructure the composition of regional economic activity along the path to net-zero emissions by 2045. Past structural changes, like the transition away from coal, do not augur well for displaced workers and regions (CEA, 2024). The evidence suggests that policymakers may need to consider expanding options in Kern through, for example, place-based policy. [^1] What will happen in California as 2045 approaches is impossible to know with certainty. In particular, a backward-looking structural VAR model will not be a good predictor of abrupt change. Instead, we view Kern’s responses to identified structural innovations as suggestive evidence that the Kern economy depends on oil. The exercise shows how Kilian’s (2009a) pioneering methods and conclusions apply more generally.
@@
Questions about how oil affects the US economy are not fully answered. Further research will benefit from other contributions of our replication. We port Kilian’s original MATLAB code to R. Some researchers may prefer to work in R as the R ecosystem makes tidying data easier. (Our code embraces the tidyverse.) We use the targets package to ensure replicability. The targets package provides Make-like tools that automatically make any necessary changes to dependencies when any change is registered. In addition, we retrieve data from the US Energy Information Administration and FRED using code, which makes estimating the most recent structural shocks as straightforward as rerunning the code. These resources will make addressing questions easier in the ongoing debate about the effects oil-price shocks have on the US economy. [^2] An introduction to Make is provided by Jesús Fernández-Villaverde in lecture notes titled “Computational Methods for Economists,” which are available at https://www.sas.upenn.edu/~jesusfv/teaching.html
We present a replication of Kilian’s (2009a) results in section 2. We conclude that Kilian (2009a) should be praised and thanked for writing and sharing code that makes reproducing results straightforward (Kilian, 2009b). Using the posted data and our code, we are able to exactly replicate Kilian’s (2009a) results.
@@
The replication comprises a description of the data (section 2), a specification of the structural VAR model (section 2.1), a presentation of the estimated impulse response functions (section 2.2), a decomposition of the real price of oil into identified components (section 2.3), and a second-stage analysis of how US real GDP and the Consumer Price Index respond to the identified structural shocks. In section 3, building upon an analysis undertaken by Kilian and Lütkepohl (2017), we strengthen Kilian’s (2009a) conclusions by estimating the structural VAR model using data from February 1974 to January 2025 and constructing confidence intervals for impulse responses that are robust to arbitrary forms of conditional heteroskedasticity. To demonstrate the wider applicability of the results, in section 4, we analyze how the global oil market affects outcomes in a local labor market.
@@ 2 Replication
Kilian’s (2009a) pioneering model of the global crude-oil market specifies three determinants. The methodology separates a single supply shock from two demand shocks. The oil-supply shock reflects unexpected disruptions to crude-oil availability. The aggregate-demand shock reflects the consensus that shifts in global real economic activity shift current demand for crude oil (Kilian and Zhou, 2018). And the precautionary-demand shock reflects how demand may shift based on uncertainty about future unavailability of oil. Alquist and Kilian (2010) make this point, among others, in an economy that receives a random endowment of oil each period. They show that an increase in the risk of future availability of oil raises current demand.
Three variables comprise Kilian’s (2009a) vector autoregressive model of the global crude-oil market: (i) the log difference of global crude-oil production, (ii) a measure of cyclical variation in global real economic activity, and (iii) the log of the real price of oil. Data on global crude-oil production are available from the US Energy Information Administration (EIA). Kilian’s (2009a) measure of cyclical variation in global real economic activity is made available by the Federal Reserve Bank of Dallas through the FRED database. Kilian (2019) reviews the index and corrects an error that was made when the original series was constructed. The real price of oil is constructed by deflating US refiners’ acquisition cost of imported crude oil (available from the EIA) by the Consumer Price Index for All Urban Consumers. Some details about the series and how we accessed them are included in an appendix in section A.1.333The appendix also reports a sensitivity analysis, in which we replace Kilian’s (2009a) index with the monthly growth rate of OECD+6 industrial production, which is an index of industrial production for OECD (Organization for Economic Co-operation and Development) countries and six non-member countries (Brazil, China, India, Indonesia, the Russian Federation, and South Africa). Further discussion is provided in the appendix.
@@ 2.1 The structural VAR model
To be precise about how the determinants of the global market for crude oil are related, consider a VAR model for $y_{t}=\left(\Delta\text{oil production},\text{real activity},\text{real price of oil}\right)^{\prime}$. The structural VAR representation is
@@
where the $B_{i}$ are $\times 3$ 3 parameter matrices and $\varepsilon_{t}\sim\left(0,\Sigma_{\varepsilon}\right)$; that is, $\varepsilon_{t}$ denotes a vector of serially and mutually uncorrelated innovations. The covariance matrix $\Sigma_{\varepsilon}$ is diagonal. The corresponding reduced-form model is
@@
Within the system of three equations, the first equation indicates that the stochastic short-run supply curve is vertical. In other words, oil production does not respond to innovations in demand within the month. The maintained assumption is that decisions about oil supply are based on medium-run expectations and, because adjusting supply is costly, short-run demand innovations are postulated to have a negligible affect on medium-run expectations. Evidently, the model allows energy prices to respond to all past information. There are two demand shocks and each is identified by delay restrictions. The short-run demand curve is downward sloping and can be shifted by innovations to aggregate demand and innovations to oil-specific demand, such as precautionary-demand innovations (Alquist and Kilian, 2010). The delay restriction is that oil-specific demand shocks do not affect global real economic activity as measured by fluctuations in shipping rates relative to trend. These postulations are further discussed by Kilian (2008) and Kilian (2009a).
@@
Inference is based on the wild bootstrap discussed by Gonçalves and Kilian (2004) with 1,000 replications. Kilian (2009a) uses 2,000 replications but the results are similar. The wild bootstrap allows for conditional heteroskedasticity of unknown form. But the asymptotic validity of the wild bootstrap is for the reduced-form impulse responses and not the asymptotic distribution of the error–covariance matrix. This feature is shown by Brüggemann, Jentsch, and Trenkler (2016) and discussed by Kilian and Lütkepohl (2017). In section 3, we follow Kilian and Lütkepohl (2017) and adopt the inference procedures proposed by Brüggemann, Jentsch, and Trenkler (2016) to investigate whether the larger confidence intervals matter in practice.
@@
The structural impulse responses define the responses of elements of $y_{t}$ to one-time impulses in $\varepsilon_{t}$. Figure 1 shows how global oil production, real economic activity, and the real price of oil respond to one-standard deviation structural innovations. The innovations are normalized so that each innovation will tend to raise the price of oil or, put differently and slightly imprecisely, figure 1 shows the effects of negative supply shocks and positive demand shocks. Figure 1 is comparable to figure 3 in Kilian (2009a).
@@
Note: Solid lines show the point estimates and the shaded regions show two-standard-error bands. Estimates are based on the structural VAR model in (1), using the original data (Kilian, 2009b) (February 1973–December 2007). Confidence intervals were constructed using a recursive-design wild bootstrap. Dotted lines show the point estimates and two-standard-error bands produced by Kilian’s (2009a) code. The striped effect indicates the two sets of point estimates coincide.
Figure 1 uses data provided by Kilian (2009b). The data start in February 1973 and end in December 2007. The goal of figure 1 is to assess whether the R code written for the replication yields similar results to the MATLAB code that produces the original graphics.
Figure 1 confirms that the two sets of code produce identical results. The solid, blue lines show the point estimates of the structural impulse responses. They match those produced by the code provided by Kilian (2009a) exactly. Kilian’s (2009a) original estimates are shown using red, dotted lines. These overlay the solid, blue lines, which produces a striped effect. The shaded areas depict the two-standard-error bands constructed by the wild bootstrap. The dotted, red lines show those produced by the original code. Again, there is overwhelming overlap, which gives confidence that the two sets of codes are working.
Figure 2 uses our code base and our data over the period that overlaps the original sample period. There are three differences to keep in mind when comparing figure 1 to figure 2. First, for the analysis in figure 2, all the data were retrieved by our code from the sources listed in the text. Second, the updated data on production start in February 1974. Third, the data on global real economic activity have been corrected (Kilian, 2019). Notwithstanding these differences, the patterns are strikingly similar.
@@
An unexpected disruption to oil supply causes an immediate decline in oil production. The decline is partially reversed within the first 5 months, and “this pattern is consistent with the view that oil supply contractions in one region tend to trigger production increases elsewhere in the world” (Kilian, 2009a, 1061–1062). The fall in production causes a rise in the price of oil, but the rise is not overwhelmingly statistically significant. In figure 2, the disruption has little effect on real economic activity, which differs slightly from Kilian’s (2009a) observation that a supply disruption “causes a small temporary reduction of real economic activity that is partially statistically significant” (1,062).
Turning to the second row of figure 2, the effect of an unanticipated rise in aggregate demand causes oil production to rise about 6 months after the shock. The expansion lasts until 15 months after the shock. The unanticipated rise in aggregate demand causes a significant rise in global real economic activity. The effect is persistent. Expansions in real economic activity cause a significant rise in the real price of oil. The real price of oil rises on impact (although not statistically significantly) and the rise continues for a year.
The third row of figure 2 shows the effects of an unanticipated rise in oil-specific demand. An unanticipated rise causes oil production to briefly fall, but over the 15-month horizon, oil-specific-demand increases do not cause oil production to change much. Real economic activity in response to the unanticipated rise in oil-specific demand expands—albeit modestly, if at all. What is most clear, though, is that the unanticipated rise in oil-specific demand causes the real price of oil to jump upward immediately and continue to rise for a few months, which is consistent with models of precautionary demand. Within this class of models, this overshooting feature is discussed by Kilian (2009a) and Alquist and Kilian (2010, figure 3b, 561). Finally, the rise in the real price of oil is highly statistically significant.
Our results are entirely consistent with the results in Kilian (2009a).
@@ 2.3 The cumulative effect of oil demand and oil supply shocks on the real price of oil
How much each structural shock explains of the historically observed price of crude oil can be assessed through a historical decomposition of the covariance-stationary VAR model. The historical decomposition decomposes $y_{t}$ into structural innovations through its moving-average representation. It attributes the cumulative effect of the structural innovations in oil production, real economic activity, and oil-specific demand to the real price of oil at each point in time (Kilian and Lütkepohl, 2017).
Figure 3 shows the cumulative contribution of oil-supply, aggregate-demand, and oil-specific-demand shocks to the real price of oil in separate panels. Each panel shows 4 lines. The top panel shows the cumulative effect of oil-supply shocks. The solid, purple line is generated from the original code and original data. The dashed, beige line is generated by using the replication code together with the original data. The two series are numerically equal, which can be seen by the striped pattern, the effect of the two series overlapping. The dot–dashed, orange line is generated by using our replication code and the updated data through December 2007. The dotted, pink line is generated by using the replication code and updated data through January 2025.
@@
Note: Top, how much oil-supply shocks explain of the historically observed fluctuations in the real price of crude oil; middle, how much aggregate-demand shocks explain of the historically observed fluctuations in the real price of crude oil; bottom, how much oil-specific-demand shocks explain of the historically observed fluctuations in the real price of crude oil. Estimates are based on the structural VAR model in (1). The solid, purple line is generated by the original code and data. The dashed, beige line, which lies on top of the purple line, is generated by the replication code and the original data. The dot–dashed, orange line is generated by the replication code using data from February 1974 to December 2007. The dotted, pink line is generated by the replication code using data through January 2025.
@@
The middle panel of figure 3 shows that aggregate-demand shocks cause prolonged and persistent changes in the price of oil. The bottom panel of 3 shows that shocks to oil-specific demand are responsible for the fluctuations in the price of oil. As Kilian (2009a) points out, this statistical understanding of crude-oil prices is inconsistent with narrative hypotheses about supply disruptions in the Middle East. There may be supply disruptions, but the disruptions affect firms that purchase oil, fearing further supply disruptions. These precautionary-demand motives cause the price of oil to fluctuate.
Figure 3 is directly comparable to Kilian’s (2009a) figure 4. The conclusions drawn from both figures agree. Therefore, going forward results will be shown for estimates produced using the full sample.
2.4 Demand and supply shocks in the global oil market affect the US economy in different ways